### v1

@@ 6.2 Viscous-box tests
Following [9], we set the initial velocity on $\Omega=[0,1]^{2}$ to
$\mathbf{u}(x,y)=\left(\begin{array}[]{c}\,\sin^{2}(\pi x)\,\sin(2\pi y)\\ \,-\sin(2\pi x)\,\sin^{2}(\pi y)\end{array}\right)$ (72)
and advance cell-averaged initial values from $t_{0}=0$ to $t_{e}=0.5$ on four successively refined uniform grids. Errors and convergence rates in the cases of $\mathrm{Re}=10^{4}$ and $\mathrm{Re}=10^{2}$ are shown in Tables 3 and 4, respectively.
Table 3: Errors and convergence rates of GePUP-ES-SDIRK for the 2D viscous-box test with Re=$10^{4}$, $t_{0}=0.0$, $t_{e}=0.5$, $\mathrm{Cr}=0.5$ and $\lambda=1$.
@@
Table 4: Errors and convergence rates of GePUP-ES-SDIRK for the 2D viscous-box test with $\text{Re}=10^{2}$, $t_{0}=0.0$, $t_{e}=0.5$, $\mathrm{Cr}=0.1$ and $\lambda=1$.
@@
(a) $\mathcal{E}_{h}^{n}$ for Re=$10^{2}$
(b) $\mathcal{E}_{h}^{n}$ for Re=$10^{4}$
(c) $|r_{h}^{n}-1|$ for Re=$10^{2}$
(d) $|r_{h}^{n}-1|$ for $\mathrm{Re}=10^{4}$
@@
(e) $\left\|\mathbf{D}\mathbf{W}_{h}^{n}\right\|_{\mathcal{C}}$ for Re=$10^{2}$
@@
(f) $\left\|\mathbf{D}\mathbf{W}_{h}^{n}\right\|_{\mathcal{C}}$ for Re=$10^{4}$
@@
Figure 4: Results of the 2D viscous-box test with $h=\frac{1}{512}$ and $\lambda=1$. The left column contains those with $\mathrm{Re}=100$ and $\mathrm{Cr}=0.1$ while the right column those with $\mathrm{Re}=10^{4}$ and $\mathrm{Cr}=0.5$. The abscissa in all subplots is time.
@@
For $\mathrm{Re}=10^{4}$, convergence rates of the velocity in all norms are close to four. In contrast, those for $\mathrm{Re}=10^{2}$ are close to four in the $L_{1}$-norm and the $L_{2}$-norm, but are around 2.5 in the $L_{\infty}$-norm. Accordingly, convergence rates of $\nabla q$ in the $L_{\infty}$-norm for $\mathrm{Re}=10^{4}$ are also substantially higher than those for $\mathrm{Re}=10^{2}$.
@@
These results are not out of expectations. We have proved in Theorem 5 that the pressure $q$ converges to the pressure $p$ and have shown in Section 2.3 that the pressure gradient $\nabla p$ in the INSE can be split into two parts $\nabla p=\nabla p_{c}+\nu\nabla p_{s}$, where $\nabla p_{s}$ responses to the Laplace-Leray commutator. When $\nu$ is sufficiently large, $\nu\nabla p_{s}$ dominates $\nabla p_{c}$ and accounts for the bulk of $\nabla p$. Cozzi and Pego [38] showed that $\|\nabla p_{s}\|$ may not be bounded at a boundary point that is not $C^{3}$. As a practical interpretation, the pressure could develop steep gradient at a $C^{1}$ discontinuity of the domain boundary for low-Reynolds-number flows. Therefore, we believe that the order reduction in the case of Re=$10^{2}$ is caused by the dominance of $\nu\nabla p_{s}$ and the sharp corners ($C^{1}$ discontinuities) of the square domain.
@@
As shown in Figures 4(a,b), the modified energy $\mathcal{E}_{h}^{n}$ decreases monotonically over the entire simulation for both $\mathrm{Re}=10^{4}$ and $\mathrm{Re}=10^{2}$. It is also clear that the modified energy decreases faster in the higher-viscosity case. This confirms Theorem 10. Figures 4(c,d) show that the deviation of SAV from 1 is at most $10^{-10}$, thus the (unmodified) kinetic energy also decreases monotonically.
@@
For both $\mathrm{Re}=10^{4}$ and $\mathrm{Re}=10^{2}$, the evolution of the $L_{2}$-norm of velocity divergence shown in Figures 4(e,f) has essentially the same pattern: the $L_{2}$-norm first increases to a local maximum and then decreases. This pattern is different from that shown in Figures 3(c,d), but can still be very well explained by (71). Since the initial velocity (72) is perfectly divergence-free, the magnitude of $\|\mathbf{G}\mathbf{D}\mathbf{W}^{(i)}_{h}\|_{\mathcal{C}}$ is small in the early simulation stage, during which the inequality (71) has no control over the discrete velocity divergence yet. However, as the discrete divergence accumulates to the point when the RHS of (71) gets dominated by its first term, the inequality (71) takes effect and forces the discrete divergence to decrease.
@@ 7 Conclusions
We have shown that the INSE with no-slip conditions can be equivalently reformulated as variants of the GePUP formulation [5], where the main evolutionary variable is a non-solenoidal velocity with electric boundary conditions whose divergence, controlled by a heat equation with homogeneous Dirichlet boundary conditions, decays exponentially. This GePUP-E reformulation is suitable for numerically solving the INSE because
@@
The next step along this line of research is to improve our current spatial discretizations so that fully discrete algorithms that ensure decays of velocity divergence and total kinetic energy can be constructed. Additionally, we are considering other types of boundary conditions such as the nonhomogeneous Dirichlet conditions, the radiation conditions, and mixed conditions. Also, we are currently augmenting the GePUP-ES solver to adaptive mesh refinement and parallel computing for an enhanced resolution of the multiple time scales and length scales in flows at moderate or high Reynolds numbers. Another work in progress is the development of GePUP solvers for the INSE with moving boundaries.
Acknowledgments
The authors thank Prof. Buyang Li from the Hong Kong Polytechnic University for helpful discussions.

### v2

@@ 6.2 Viscous-box tests
Following bell89:_secon_order_projec_method_incom , we set the initial velocity on $\Omega=[0,1]^{2}$ to
$\mathbf{u}_{0}(x,y)=\left(\begin{array}[]{c}\,\sin^{2}(\pi x)\,\sin(2\pi y)\\ \,-\sin(2\pi x)\,\sin^{2}(\pi y)\end{array}\right)$ (67)
and advance cell-averaged initial values from $t_{0}=0$ to $t_{e}=0.5$ on four successively refined uniform grids. The Courant number is defined as $\mathrm{Cr}:=\frac{k}{h}\|\mathbf{u}_{0}\|_{\infty}$, where $\|\mathbf{u}_{0}\|_{\infty}$ is the max-norm of the initial velocity $\mathbf{u}_{0}$.
Table 3: Errors and convergence rates of the GES scheme in Table 1 for solving the 2D viscous-box test with Re $10^{4}$, $t_{0}=0.0$, $t_{e}=0.5$, $\mathrm{Cr}=0.5$ and $\lambda=1$.
@@
Table 4: Errors and convergence rates of the GES scheme in Table 1 for solving the 2D viscous-box test with $\text{Re}=10^{2}$, $t_{0}=0.0$, $t_{e}=0.5$, $\mathrm{Cr}=0.1$ and $\lambda=1$.
@@
(a) $\mathcal{E}_{h}^{n}$ for Re $10^{2}$ and $\mathrm{Cr}=0.1$
(b) $\mathcal{E}_{h}^{n}$ for Re $10^{4}$ and $\mathrm{Cr}=0.5$
(c) $|r_{h}^{n}-1|$ for Re $10^{2}$ and $\mathrm{Cr}=0.1$
(d) $|r_{h}^{n}-1|$ for $\mathrm{Re}=10^{4}$ and $\mathrm{Cr}=0.5$
@@
(e) $\left\|\mathbf{D}\mathbf{W}_{h}^{n}\right\|_{\mathcal{C}}$ (solid markers) and $\left\|\mathbf{D}\mathbf{U}_{h}^{n}\right\|_{\mathcal{C}}$ (hollow markers) for Re $10^{2}$ and $\mathrm{Cr}=0.1$
@@
(f) $\left\|\mathbf{D}\mathbf{W}_{h}^{n}\right\|_{\mathcal{C}}$ (solid markers) and $\left\|\mathbf{D}\mathbf{U}_{h}^{n}\right\|_{\mathcal{C}}$ (hollow markers) for Re $10^{4}$ and $\mathrm{Cr}=0.5$
@@
Figure 4: Results of the GES scheme in Table 1 for solving the 2D viscous-box test with $h=\frac{1}{512}$ and $\lambda=1$; see (64) and (65) for precise definitions of $\left\|\cdot\right\|_{\mathcal{C}}$ and $\mathcal{E}_{h}^{n}$. The abscissa in all subplots is time.
@@
Errors and convergence rates in the cases of $\mathrm{Re}=10^{4}$ and $\mathrm{Re}=10^{2}$ are shown in Tables 3 and 4, respectively. For $\mathrm{Re}=10^{4}$, convergence rates of the velocity in all norms are close to four. In contrast, those for $\mathrm{Re}=10^{2}$ are close to four in the $L_{1}$-norm and the $L_{2}$-norm, but are around 2.5 in the $L_{\infty}$-norm. Accordingly, convergence rates of $\nabla q$ in the $L_{\infty}$-norm for $\mathrm{Re}=10^{4}$ are also substantially higher than those for $\mathrm{Re}=10^{2}$.
@@
These results are not out of expectations. We have proved in Theorem 3.7 that the pressure $q$ converges to the pressure $p$ and have shown in Section 2.3 that the pressure gradient $\nabla p$ in the INSE can be split into two parts $\nabla p=\nabla p_{c}+\nu\nabla p_{s}$, where $\nabla p_{s}$ responses to the Laplace-Leray commutator. When $\nu$ is sufficiently large, $\nu\nabla p_{s}$ dominates $\nabla p_{c}$ and accounts for the bulk of $\nabla p$. Cozzi and Pego cozzi11:_laplac_leray showed that $\|\nabla p_{s}\|$ may not be bounded at a boundary point that is not $C^{3}$. As a practical interpretation, the pressure could develop steep gradient at a $C^{1}$ discontinuity of the domain boundary for low-Reynolds-number flows. Therefore, we believe that the order reduction in the case of Re $10^{2}$ is caused by the dominance of $\nu\nabla p_{s}$ and the sharp corners ($C^{1}$ discontinuities) of the square domain.
@@
As shown in Figure 4(a,b), the modified energy $\mathcal{E}_{h}^{n}$ decreases monotonically over the entire simulation for both $\mathrm{Re}=10^{4}$ and $\mathrm{Re}=10^{2}$. It is also clear that $\mathcal{E}_{h}^{n}$ decreases faster in the higher-viscosity case of Re $10^{2}$, which confirms Theorem 5.4. Figure 4(c,d) show that the deviation of SAV from 1 is at most $10^{-10}$, thus the (unmodified) kinetic energy also decreases monotonically.
@@
For both $\mathrm{Re}=10^{4}$ and $\mathrm{Re}=10^{2}$, the evolution of the $L_{2}$-norm of velocity divergence shown in Figure 4(e,f) has essentially the same pattern: the $L_{2}$-norm first increases to a local maximum and then decreases. This pattern is different from that shown in Figure 3(c,d), but can still be very well explained by (66). Since the initial velocity (67) is divergence-free, the magnitude of $\|\mathbf{G}\mathbf{D}\mathbf{W}^{(i)}_{h}\|_{\mathcal{C}}$ is small in the early simulation stage, during which the inequality (66) has no control over the discrete velocity divergence yet. However, as the discrete divergence accumulates to the point when the RHS of (66) gets dominated by its first term, the inequality (66) takes effect and forces the discrete divergence to decrease.
6.3 Viscous-box tests with initially non-solenoidal velocity
Tests in this subsection are the same as those in Section 6.2 except that the initial condition is set to the following non-solenoidal velocity,
$\begin{array}[]{l}\mathbf{w}_{0}=\mathbf{u}_{0}+\epsilon\nabla\phi,\quad\text{ with }\phi=\frac{1}{2\pi^{2}}\sin^{2}(\pi x)\sin^{2}(\pi y),\end{array}$ (68)
where $\mathbf{u}_{0}$ is the initially solenoidal velocity in (67), $\epsilon\nabla\phi$ the perturbation of $\mathbf{u}_{0}$ away from the divergence-free space, and $\epsilon>0$ the parameter to control the magnitude of the perturbation.
(a) $\epsilon=1$, $t=t_{0}=0$
(b) $\epsilon=0$, $t=t_{0}=0$
(c) $\epsilon\!=\!10^{-3}$, $t=t_{0}=0$
(d) $\epsilon=1$, $t=0.5$, Re $10^{4}$
(e) $\epsilon=0$, $t=0.5$, Re $10^{4}$
(f) $\epsilon\!=\!10^{-3}$, t=0.5t\!=\!0.5, Re $10^{4}$
(g) $\epsilon=1$, $t=0.5$, Re $10^{2}$
(h) $\epsilon=0$, $t=0.5$, Re $10^{2}$
(i) $\epsilon\!=\!10^{-3}$, $t=0.5$, Re $10^{2}$
[…]
Hereafter we denote by $\mathbf{U}^{n}_{\epsilon=0}$ the reference solution obtained from the divergence-free initial velocity $\mathbf{u}_{0}$, i.e., $\mathbf{U}^{n}_{h}$ in Section 6.2, and denote by $\mathbf{U}^{n}_{\epsilon>0}$ the perturbed solution computed from a perturbed initial velocity $\mathbf{w}_{0}$.
Streamlines of the initial velocity and the final solutions are plotted in Figure 5, where no qualitative differences can be observed between corresponding plots in the third and the second columns, due to the small perturbation magnitude $\epsilon=10^{-3}$. In contrast, for $\epsilon=1$, the perturbation is sufficient to generate different patterns for the initial velocity, cf. Figure 5(a,b). For example, the streamlines in Figure 5(a) are not closed due to the large magnitude of velocity divergence. However, the perturbation effects have already died out at $t=0.5$ and the final perturbed solution $\mathbf{U}^{n}_{\epsilon=1}$ is visually indistinguishable from the unperturbed solution $\mathbf{U}^{n}_{\epsilon=0}$ for both Re $10^{4}$ and Re $10^{2}$; see Figure 5(d,e) and Figure 5(g,h).
How fast does the perturbed solution $\mathbf{U}^{n}_{\epsilon>0}$ converge to the solution $\mathbf{U}^{n}_{\epsilon=0}$? The answer lies in Figure 6, where the deviation of $\mathbf{U}^{n}_{\epsilon>0}$ from $\mathbf{U}^{n}_{\epsilon=0}$ decays by a factor of at least $10^{3}$ within the first time step for all test cases. Furthermore, although the norm $\|\mathbf{U}^{n}_{\epsilon>0}-\mathbf{U}^{n}_{\epsilon=0}\|_{\mathcal{C}}$ appears to be proportional to the perturbation magnitude, the reduction rate during the first time step is largely independent of the time step size; see Figure 6(b,c) and Figure 6(e,f).
(a) Re $10^{2}$, Cr =0.1=0.1
(b) Re $10^{2}$, Cr =0.1=0.1
(c) Re $10^{2}$, Cr =0.01=0.01
(d) Re $10^{4}$, Cr =0.5=0.5
(e) Re $10^{4}$, Cr =0.5=0.5
(f) Re $10^{4}$, Cr =0.05=0.05
Figure 6: Evolutions of $\|\mathbf{U}^{n}_{\epsilon>0}-\mathbf{U}^{n}_{\epsilon=0}\|_{\mathcal{C}}$, i.e., the $L_{2}$-norm (64) of the difference between the computed velocities of the perturbed viscous-box tests in Section 6.3 on a uniform spatial grid $h=\frac{1}{512}$. The markers $\bullet$,” $\blacksquare$,” $\blacktriangle$” and “++” represent values of $\|\mathbf{U}^{n}_{\epsilon>0}-\mathbf{U}^{n}_{\epsilon=0}\|_{\mathcal{C}}$ for $\epsilon=1$, $0.1$, $10^{-2}$, and $10^{-3}$, respectively. The abscissas in subplots (b,c,e,f) and (a,d) represent the index of time steps and the simultation time, respectively.
(a) $\mathcal{E}^{n}_{h}$ for Re $10^{2}$ and Cr = 0.1
(b) $\mathcal{E}^{n}_{h}$ for Re $10^{4}$ and Cr = 0.5
(c) $|r^{n}_{h}-1|$ for Re $10^{2}$ and Cr = 0.1
(d) $|r^{n}_{h}\!-\!1|$ for Re $10^{4}$ and Cr = 0.5
(e) $\|\mathbf{DU}^{n}_{\epsilon}\|_{\mathcal{C}}$ for Re $10^{2}$ and Cr = 0.1
(f) $\|\mathbf{DU}^{n}_{\epsilon}\|_{\mathcal{C}}$ for Re $10^{4}$ and Cr = 0.5
Figure 7: Evolutions of $\mathcal{E}^{n}_{h}$, $|r^{n}_{h}-1|$, and $\|\mathbf{DU}^{n}_{\epsilon}\|_{\mathcal{C}}$ produced by the GES scheme in Table 1 for solving the perturbed viscous box tests on a uniform grid of $h=\frac{1}{512}$; see (64) and (65) for precise definitions of $\left\|\cdot\right\|_{\mathcal{C}}$ and $\mathcal{E}_{h}^{n}$. The dashed curves represent results with $\epsilon=0$ (i.e., those in Figure 4) while those marked by $\bullet$,” $\blacksquare$,” $\blacktriangle$,” and $\bigstar$” correspond to $\epsilon=1$, $10^{-1}$, $10^{-2}$ and $10^{-3}$, respectively. The abscissa of all subplots is time.
For two curves very close to each other, their tangent vectors might differ largely, so do other geometric quantities such as the curvature. Likewise, $\|\mathbf{U}^{n}_{\epsilon>0}-\mathbf{U}^{n}_{\epsilon=0}\|_{\mathcal{C}}$ being small does not imply $\|{\mathbf{D}}\mathbf{U}^{n}_{\epsilon>0}-{\mathbf{D}}\mathbf{U}^{n}_{\epsilon=0}\|_{\mathcal{C}}$ being small. This issue is addressed by Figure 7, which shows the perturbation effects on evolutions of the modified energy, the SAV, and the velocity divergence.
Compared to Figure 4(a,b), Figure 7(a,b) feature a decrease of the modified energy during the very first time step for $\epsilon=1$, indicating that most of the extra energy corresponding to the non-solenoidal perturbation has been removed. Values of $|r_{h}^{n}-1|$ in Figure 7(c,d) are much larger than those in Figure 4(c,d), which is still acceptable because the deviation of $r_{h}^{n}$ from 1 is about $10^{-5}$ even for the case of the largest perburbation $\epsilon=1$. In Figure 7(e,f), the evolution curve of $\|\mathbf{DU}^{n}_{\epsilon}\|_{\mathcal{C}}$ for each $\epsilon>0$ overlaps with that of $\epsilon=0$ except for the first four steps.
(a) Re $10^{2}$ and Cr = 0.1
(b) Re $10^{4}$ and Cr = 0.5
(c) Re $10^{2}$ and Cr = 0.01
(d) Re $10^{4}$ and Cr = 0.05
Figure 8: Evolution of discrete velocity divergence produced by the GES scheme in Table 1 for solving the perturbed viscous box tests on a uniform grid of $h=\frac{1}{512}$. The abscissa of all subplots is the index of time steps. Subplots (a,b) can be considered as more detailed versions of Figure 7 (e,f). Values of $\|\mathbf{DU}^{n}_{\epsilon}\|_{\mathcal{C}}$ for $\epsilon=0,10^{-3},1$ are represented respectively by solid markers $\blacktriangle$,” $\blacksquare$,” and $\bullet$” while $\|\mathbf{DW}^{n}_{\epsilon}\|_{\mathcal{C}}$ respectively by the corresponding the hollow markers $\triangle$,” $\square$,” and $\circ$.” Results of $\epsilon=10^{-1},10^{-2}$ are in between those of $\epsilon=10^{-3},1$ and have the same qualitative pattern.
To examine what happens in these four steps, the temporal variations of the $L_{2}$-norm of the velocity divergence are plotted in Figure 8 for the two representative cases of $\epsilon=1,10^{-3}$. Different from previous figures on velocity divergence, Figure 8 contains values of both $\|\mathbf{DW}^{n}_{\epsilon}\|_{\mathcal{C}}$ and $\|\mathbf{DU}^{n}_{\epsilon}\|_{\mathcal{C}}$, which are connected by the discrete projection that approximates $\mathscr{P}$} in (48). As illustrated in Figure 8, the discretization error might increase $\|\mathbf{DW}^{n}_{\epsilon}\|_{\mathcal{C}}$ during a single time step, but the magnitude of this increase approaches a constant value after about eight time steps; furthermore, the discrete projection at the end of a time step counteracts this increase. In each subplot of Figure 8, the sequence of solid markers $\bullet$” and $\blacksquare$” during the first four time steps clearly demonstrates the fast decay of $\|\mathbf{DU}^{n}_{\epsilon}\|_{\mathcal{C}}$. In addition, a comparison of the two rows of Figure 8 shows that the adverse effects of the perturbation upon the divergence-free condition can be further reduced by decreasing the size of the initial time steps.
We have demonstrated two aspects of the prominent advantage of GePUP-ES in handling an initially non-solenoidal velocity. First, the velocity divergence decays exponentially. Second, Figures 6 and 8 suggest that the time period of adverse effects caused by the initially non-solenoidal velocity can be very much shortened by reducing the size of the first several time steps. We emphasize that theoretically there is no guarantee of the decay of velocity divergence for any initially non-solenoidal velocity yet, as the well-posedness of INSE is still one of the unsolved Millennium Prize Problems.
(a) $\mathcal{E}^{n}_{h}$
(b) $\|\mathbf{DU}^{n}_{\epsilon}\|_{\mathcal{C}}$
Figure 9: Evolution of $\mathcal{E}^{n}_{h}$ and $\|\mathbf{DU}^{n}_{\epsilon}\|_{\mathcal{C}}$ produced by applying the MOL approach to the strong form of UPPE in (13) for Re = $10^{4}$ , Cr = 0.5, and $h=\frac{1}{512}$. The curves marked by $\bullet$” and $\blacksquare$” correspond to $\epsilon=1$ and $10^{-3}$, respectively, while the dashed curve corresponds to $\epsilon=0$. The abscissa in all subplots is time.
The finite-volume based MOL scheme is also applied to the strong form of UPPE in (13) to obtain solutions of the perturbed viscous-box tests. As shown in Figure 9, both the velocity divergence and the total kinetic energy blow up, even for the case of an initially solenoidal velocity. It is also clear that the blow-up time becomes sooner as $\epsilon$ gets larger.
Results of viscous-box tests solved by GePUP, as shown in (zhang16:_GePUP, , Table 3), are very close to those in Table 3 for $\mathrm{Re}=10^{4}$. In the case of $\mathrm{Re}=10^{2}$, the error norms of $\mathbf{u}$ produced by GePUP-ES on the finest grid are approximately 20% smaller than those by GePUP, cf. Table 4 and (zhang16:_GePUP, , Table 4). For the perturbed viscous box tests in this subsection, results of GePUP also show fast decays of discrete velocity divergence. However, GePUP-ES is a decisive advance from GePUP in that, in the semi-discrete case, the decay of velocity divergence and kinetic energy can be rigorously proven.
@@ 7 Conclusions
We have shown that the INSE with no-slip conditions can be equivalently reformulated as variants of the GePUP formulation zhang16:_GePUP , where the main evolutionary variable is a non-solenoidal velocity with electric boundary conditions whose divergence, controlled by a heat equation with homogeneous Dirichlet boundary conditions, decays exponentially. This GePUP-E reformulation is suitable for numerically solving the INSE because
@@
We are currently augmenting the GePUP-ES solver to parallel computing and adaptive mesh refinement zhao2025FourthOrderAMR for an enhanced resolution of the multiple time scales and length scales in flows at moderate or high Reynolds numbers. Another work in progress is the development of GePUP solvers for the INSE with irregular and moving boundaries via poised lattice generation zhang2024PLG .
The next step along this line of research is the design of fully discrete algorithms that ensure decays of velocity divergence and total kinetic energy. We also plan to consider other types of boundary conditions such as the nonhomogeneous Dirichlet conditions, the radiation conditions, and mixed conditions.
Acknowledgements.
This work was supported by the grant 12272346 from the National Natural Science Foundation of China. We acknowledge helpful discussions with Prof. Buyang Li at the Hong Kong Polytechic University. We are also grateful to one anonymous referee, whose insightful comments and suggestions lead to an improvement of this paper.
Declarations
Data Availability
The datasets generated during and/or analysed during the current study are available from the corresponding author on reasonable request.
Conflict of interest
The authors declare that they have no conflict of interest.